Calculate the enthalpy of reaction using bond enthalpies for the following reaction: Cl₂(g) + 2H₂O(g) → 4HCl(g) + O₂(g)
1
Expert's answer
2020-04-28T12:20:52-0400

Bond enthalpy of "Cl-Cl =243KJ\/mol"

Bond enthalpy of "O-H=460KJ\/mol"

Bond enthalpy of "H-Cl=432KJ\/mol"

Bond enthalpy of "O-O=498KJ\/mol"

The reaction is:

"Cl_2(g) + 2H_2O(g) \u2192 4HCl(g) + O_2(g)"

Ethalpy of reaction"=" Total Bond energy of reactants"-" Total Bond energy of products"=(243+2\times(2\times 460)-4\times 432-498)=-183KJ/mol"

Negative sign represents it is an exothermic reaction.